# The origin of supermassive black holes at cosmic dawn

### v2

@@ 5 Discussions and Conclusions
We have studied a spherically symmetric accretion flow into the combined potential of a black hole at the centre of a dark matter halo. We have considered both a Hernquist and an NFW potential for the halo. The accretion flow exhibits two critical points without a central black hole. The trans-sonic flow passes through the outer critical point, then circles around the inner critical point, and does not proceed all the way to the centre.
@@
The flow has three critical points in the presence of a black hole. The trans-sonic solution is possible either through the innermost or the outermost critical point. The two solutions operate exclusively in two separate regimes of $\alpha$, roughly the ratio of the halo velocity dispersion and the sound speed. For a constant temperature, $\alpha$ decreases with redshift. Initially, due to higher value of $\alpha$, the trans-sonic flow proceeds through the innermost critical point with a lower accretion rate. Subsequently as $\alpha$ decreases below a transition value, the flow proceeds through the outermost critical point at a relatively high and constant dimensionless accretion rate.
@@
We provide the general expressions for the mass accretion rates in the combined potential of a black hole and a dark matter halo, for both a Hernquist (Eq. 15) and an NFW halo (Eq. 25). The mass accretion rate has two distinct phases, depending on whether the trans-sonic solution proceeds through the innermost critical point (due to the black hole) or the outermost critical point (due to the halo). The two phases of accretion for the SMBH are proposed in simulation studies as well, albeit for different reasons that connect SMBH growth to galaxy formation (Dubois et al., 2015; Anglés-Alcázar et al., 2017) and mass reservoir in the halo (e.g. Bower et al., 2017; McAlpine et al., 2018). Moreover, the seed black hole mass used in the simulations is much higher ($\sim 10^{5}$ M) than the stellar mass seed considered in this paper. We emphasise that our finding of the two phases of accretion is based purely on hydrodynamic grounds, and of fundamental nature.
@@
In the initial phase, the mass accretion rate is due to the black hole, and it is low. After the initial phase, the mass accretion rate increases sharply (at redshift $\approx 12.5$ for a halo of present-day mass $10^{13}$ M). The growth of black holes to supermassive scales occurs during this phase (Fig. 6 top panel).
@@
Although the episode of super-Eddington is brief but it may hinder accretion through increased radiative feedback. However, during the short super-Eddington period, the black hole mass is low and likely insufficient (Fig. 6 top panel) to curtail accretion through radiation feedback and outflows (Silk & Rees, 1998; King, 2003; Li, 2012).
Our solution is for an isothermal case, whereas the exact thermodynamics of the flow is more involved, with cooling and shock heating, that needs to be investigated in detail. Nevertheless, our findings are a fundamental character of the hydrodynamic accretion likely to re-appear in a general polytropic case or a case with cooling and heating.
The actual accretion in the haloes is likely filamentary, as often reported in cosmological simulations, with further virialization and gas cooling in haloes. Also, a part of the cosmological infall leads to the development of galaxy and its spheroidal . Nonetheless, the mass accretion rate we provide (Eq. 26) is relevant for the sub-grid modelling of accretion and growth of SMBHs in cosmological simulations, as has been the standard result of Bondi (Bondi, 1952). Evolution of the black hole and the in-situ development of galactic components need investigation in the context of two hydrodynamic phases of accretion flow reported in this study.
We conclude that the stellar mass seed black holes in haloes grow via two phases of trans-sonic accretion, beginning with a phase of a low accretion rate and negligible growth due to the trans-sonic flow through the innermost critical point; followed by a phase of a high accretion rate due to the trans-sonic flow through the outermost critical point. The accretion is briefly super-Eddington, for roughly $100$ Myr, during the transition between the two phases. During the short super-Eddington episode, the accretion is unlikely to stop as the black hole mass is low to have a significant radiative feedback. Subsequently in the phase of rapid growth, the accretion becomes sub-Eddington as the black hole grows to supermassive scales. The growth history of black holes reported in this paper explains the origin of the SMBHs detected at $z\geq 6$ by the JWST (Larson et al., 2023; Maiolino et al., 2023a, b; Kokorev et al., 2023) and quasars observed at $z\geq 6$ (Fan et al., 2001, 2003; Mortlock et al., 2011; Bañados et al., 2016, 2018; Reed et al., 2019; Yang et al., 2020; Wang et al., 2021; Matsuoka et al., 2018b, a, 2019a; Inayoshi et al., 2020).
@@ Acknowledgements
MS thanks the Department of Science and Technology (DST) and Science and Engineering Research Board (SERB) India, for supporting the research through grants SRG/2022/001137 and MTR/2022/000548. RS is grateful to the ministry of education (MoE) and DST India for supporting this work. We thank an anonymous referee for constructive comments that improved the manuscript.
@@
The data underlying this article will be shared on reasonable request to the corresponding author.
@@ Appendix A Derivation of transition $\alpha$ that separates two phases of accretion
For the trans-sonic solution passing through the outermost critical point, $\mathcal{M}=1$ and $x\approx 0.5$, we obtain $C=C_{\rm T}$ where
$C_{\rm T}\approx 2\ln{2}-3/2\ .$ (29)
At $\alpha=\alpha_{\rm T}$ this solution also passes through the innermost critical point, for which we can use $\mathcal{M}=1$, $x\approx\beta/2$ and $C=C_{\rm T}$ in Eq. 28 to obtain,
@@
The expressions for $\alpha_{\rm T}$ can also be derived by equating the mass accretion rates in two asymptotic regimes. For example, by equating the mass accretion rates in equations 16 and 18 for the Hernquist halo, and by equating the mass accretion rates in Eq. 26 for the NFW halo.

### v3

@@ 5 Summary and Conclusions
We have studied a spherically symmetric accretion flow into the combined potential of a black hole at the centre of a dark matter halo. We have considered both a Hernquist and an NFW potential for the halo. The accretion flow exhibits two critical points without a central black hole. The trans-sonic flow passes through the outer critical point, then loops around the inner critical point, and does not proceed all the way to the centre.
@@
The flow has three critical points in the presence of a black hole. The trans-sonic solution is possible through the innermost or the outermost critical point. The two solutions operate exclusively in two separate regimes of $\alpha$, roughly the ratio of the halo velocity dispersion and the sound speed. For a constant temperature, $\alpha$ decreases with redshift. Initially, due to a higher value of $\alpha$, the trans-sonic flow proceeds through the innermost critical point with a lower accretion rate. Subsequently, as $\alpha$ decreases below a transition value, the flow proceeds through the outermost critical point at a relatively high and constant dimensionless accretion rate.
@@
We provide the general expressions for the mass accretion rates in the combined potential of a black hole and a dark matter halo for both a Hernquist (Eq. 15) and an NFW halo (Eq. 25). The mass accretion rate has two distinct phases, depending on whether the trans-sonic solution proceeds through the innermost critical point (due to the black hole) or the outermost critical point (due to the halo). The two phases of accretion for the SMBH are proposed in simulation studies as well, albeit for different reasons that connect SMBH growth to galaxy formation (Dubois et al., 2015; Anglés-Alcázar et al., 2017) and mass reservoir in the halo (e.g. Bower et al., 2017; McAlpine et al., 2018). Moreover, the seed black hole mass used in the simulations is much higher ($\sim 10^{5}$ ${}_{\odot}$) than the stellar mass seed considered in this paper. We emphasize that our finding of the two phases of accretion is based on hydrodynamic grounds and of a fundamental nature.
@@
In the initial phase, the mass accretion rate is due to the black hole, and it is low. After the initial phase, the mass accretion rate increases sharply (at redshift $\approx 12.5$ for a halo of present-day mass $10^{13}$ ${}_{\odot}$). The growth of black holes to supermassive scales occurs during this phase (Fig. 6 top panel).
@@
Although the super-Eddington episode is brief, it may hinder accretion through increased radiative feedback. However, during the short super-Eddington period, the black hole mass is low and likely insufficient (Fig. 6 top panel) to curtail accretion through radiation feedback and outflows (Silk & Rees, 1998; King, 2003; Li, 2012).
Our solution is for an isothermal case, whereas the exact thermodynamics of the flow is likely more involved, with cooling and shock heating, which require detailed investigation. Nevertheless, our results highlight a fundamental characteristic of hydrodynamic accretion that is likely to re-appear in a general polytropic case or in a case with cooling and heating.
The actual accretion in the haloes is likely filamentary, as often reported in cosmological simulations, with further virialization and gas cooling effects. Also, a part of the cosmological infall leads to the development of the galaxy and its spheroidal. Nonetheless, the mass accretion rate we provide (Eq. 26) is relevant for the sub-grid modelling of accretion and growth of SMBHs in cosmological simulations, as has been the standard result of Bondi (Bondi, 1952). The evolution of the black hole and the in-situ development of galactic components need investigation in the context of two hydrodynamic phases of accretion flow reported in this study.
We conclude that the stellar mass seed black holes in haloes grow via two phases of trans-sonic accretion, beginning with a phase of a low accretion rate and negligible growth due to the trans-sonic flow through the innermost critical point, followed by a phase of a high accretion rate due to the trans-sonic flow through the outermost critical point. The accretion is briefly super-Eddington, for roughly $100$ Myr, during the transition between the two phases. During the short super-Eddington episode, the accretion is unlikely to stop as the black hole mass is too low to have significant radiative feedback. Subsequently, in the phase of rapid growth, the accretion becomes sub-Eddington as the black hole grows to supermassive scales. The growth history of black holes reported in this paper explains the origin of the SMBHs detected at $z\geq 6$ by the JWST (Larson et al., 2023; Maiolino et al., 2023a, b; Kokorev et al., 2023) and quasars observed at $z\geq 6$ (Fan et al., 2001, 2003; Mortlock et al., 2011; Bañados et al., 2016, 2018; Reed et al., 2019; Matsuoka et al., 2018b, a, 2019a; Yang et al., 2020; Wang et al., 2021; Inayoshi et al., 2020).
@@ Acknowledgements
We thank an anonymous referee for constructive comments that improved the manuscript. MS thanks the Department of Science and Technology (DST) and Science and Engineering Research Board (SERB) India for supporting the research through grants SRG/2022/001137 and MTR/2022/000548. RS is grateful to the Ministry of Education (MoE) and DST India for supporting this work.
@@
The calculations and data underlying this article will be shared upon reasonable request to the corresponding author.
@@ Appendix A Derivation of transition $\alpha$ that separates two phases of accretion
For the trans-sonic solution passing through the outermost critical point, $\mathcal{M}=1$, $x\approx 0.5$, and as $\alpha\ll 0.5$ and $\beta\ll 1$, Eq. 28 yields, $C\approx C_{\rm T}$, where,
$C_{\rm T}=2\ln{2}-3/2\ .$ (29)
At $\alpha=\alpha_{\rm T}$, the same trans-sonic solution also passes through the innermost critical point, for which we can substitute $\mathcal{M}=1$, $x\approx\beta/2$ and $C=C_{\rm T}$ in Eq. 28 to obtain,
@@
The expressions for $\alpha_{\rm T}$ can also be derived by equating the mass accretion rates in two asymptotic regimes; by equating the mass accretion rates in equations 16 and 18 for the Hernquist halo, and by equating the mass accretion rates in Eq. 26 for the NFW halo.